# Convexified Message-Passing Graph Neural Networks

### v1

@@ 4 Convexification of GCNs
As $\mathcal{F}_{\text{gcn}}$ in (3) is a non-convex set and the activation function is nonlinear, the ERM problem on GCNs is non-convex, raising the challenges mentioned in Section 1. To overcome them, we propose a procedure for making both the ERM’s predictions domain and loss function convex. Following Zhang et al. [59], we develop the class of Convexified GCNs (CGCNs). In Section 4.1, we first depict our method for linear activation functions. Given that the linear activation function would collapse all layers into one, this case is of no interest in practice. However, it supplies us with some insight into the general kernelization scheme of GCNs. In the nonlinear case (Section 4.2), we reduce the problem to the linear case by proposing a relaxation to a suitably chosen Reproducing Kernel Hilbert Space (RKHS), enabling us to obtain a convex formulation of the associated ERM problem. We then present an algorithm for learning CGCNs (Section 4.3), discuss its scalability and time complexity (Section 4.4), and establish its generalization guarantees (Section 4.5).
4.1 Convexificaition of GCNs with Linear Activation Functions
To provide the intuition for our method, we employ the nuclear norm in the simple case of the linear activation function $\sigma(\mathbf{w})=\mathbf{w}$, which provides us with a convex function class, as opposed to (3). In this case, the output of layer $\ell$ is simply $\mathcal{X}_{\ell}=\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}$ due to (1)-(2). Hence, $\mathcal{X}_{\ell}$ linearly depends on the filters $\{\mathcal{A}_{\ell k}\}_{k=0}^{K}$ of layer $\ell$, i.e., the convolution function $\Psi^{\mathcal{A}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}}\rightarrow\mathbb{R}^{n\times F_{\ell}}$ in (1) linearly depends on the parameters.
@@ 4.1 Convexificaition of GCNs with Linear Activation Functions
Nuclear Norm Relaxation. However, filters that satisfy the constraints (5a) and (5b) form a non-convex set. A standard convex relaxation of a rank constraint involves replacing it with a nuclear norm constraint. That is, controlling the sum of singular values instead of the number of non-zero ones, similarly to how the $\ell_{1}$-c can be viewed as a convex relaxation of the $\ell_{0}$-norm constraint.
Upper Bounding the Nuclear-Norm. In Appendix A, we prove that:
@@
4.2 Convexification of GCNs with Nonlinear Activation Functions
For certain nonlinear activation functions $\sigma$ and suitably chosen kernel function $\kappa$ such as the Gaussian RBF kernel, we next prove how the class $\mathcal{F}_{\text{gcn}}$ of GCNs in (3) can be relaxed to a Reproducing Kernel Hilbert Space (RKHS), reducing the problem to the linear activation case (Readers may refer to [39] for a brief overview on RKHS). Beforehand, we rephrase the filter at layer $\ell$ from (1)-(2) as follows. Denoting the $f^{\text{th}}$ column of the filters $\mathcal{A}_{\ell k}$ as $\mathbf{a}^{f}_{\ell k}$, in Appendix C we prove that:
@@ 4.2 Convexification of GCNs with Nonlinear Activation Functions
A GCN induces a nonlinear filter $\tau_{\ell}^{f}$ at each layer $\ell$ for the $f^{\text{th}}$ feature given as follows (where the activation function $\sigma$ is applied entrywise):
@@
For any graph signal $\mathcal{S}^{k}\mathcal{X}$, let $\mathbf{z}_{i}^{k}(\mathcal{X}):=[\mathcal{S}^{k}\mathcal{X}]_{i}$ be the $i^{\text{th}}$ row of $\mathcal{S}^{k}\mathcal{X}$, which is the information aggregated at node $i$ from $k$-hops away. Defining the matrix $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is $[\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k+1}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$, the output at layer $\ell$ is thus given by $\tau_{\ell}^{k}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$.
Kernelization of GCNs. In the sequel, we consider a training data $\mathcal{T}=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$, some layer of index $\ell$ and a positive semi-definite kernel function $\kappa_{\ell}:\mathbb{R}^{F_{\ell-1}}\times\mathbb{R}^{F_{\ell-1}}\rightarrow\mathbb{R}$ for layer $\ell$. Thus, for brevity, we omit indices involving $\ell$ and $\mathcal{X}_{\ell-1}^{(j)}$ when no ambiguity arises. Namely, we denote, e.g., $\mathcal{Z}_{i,j}:=\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}^{(j)})$ and $\mathbf{z}_{i,j}^{k}:=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})$.
First, we prove in Lemma 4.3 that the filter specified by (7) resides in the RKHS induced by the kernel (See Appendix D for a proof, which adapts the one by Zhang et al. [59] to GCNs). Zhang et al. [59] prove that convolutional neural networks can be contained within the RKHS induced by properly chosen kernel and activation functions. For certain choices of kernels (e.g., the Gaussian RBF kernel) and a sufficiently smooth $\sigma$ (e.g., smoothed hinge loss), we can similarly prove that the filter $\tau_{\ell}^{f}$ in (7) is contained within the RKHS induced by the kernel function $\kappa_{\ell}$. See Appendices B.1-B.2 for possible choices of kernels and Appendix B.3 for valid activation functions, as well as an elaboration on sufficient smoothness in Remark B.2.
@@
The filter $\tau^{f}_{\ell}$ at layer $\ell$ for the $f^{\text{th}}$ feature can be reparametrized as a linear combination of kernel products, reducing the problem to the linear activations case. Formally:
@@
To avoid re-deriving all results in the kernelized setting, we reduce the problem to the linear case in Section 4.1. Thus, consider the (symmetric) kernel matrix $\mathcal{K}^{k}_{\ell}\in\mathbb{R}^{nm\times nm}$, whose entry at row $(i,j)$ and column $(i^{\prime},j^{\prime})$ equals to $\kappa_{\ell}(\mathbf{z}_{i,j}^{k},\mathbf{z}_{i^{\prime},j^{\prime}}^{k})$. We factorize the kernel matrix as $\mathcal{K}^{k}_{\ell}=Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$, where $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ (as in [59], we can use, e.g., Cholesky factorization [8] with $P=nm$).
@@
Thus, the reparameterization in (8) of the filter at layer $\ell$ for the $f^{\text{th}}$ feature can be rephrased as:
$\tau^{f}_{\ell}(\mathcal{Z}_{i,j})=\sum_{k=0}^{K}\langle\mathbf{q}_{i,j}^{k},\hat{\mathbf{a}}_{\ell k}^{f}\rangle;\quad\hat{\mathbf{a}}_{\ell k}^{f}:=\sum_{(i^{\prime},j^{\prime},k)\in[n]\times[m]\times([K]\cup\{0\})}\alpha_{k,(i^{\prime},j^{\prime})}^{f}\mathbf{q}_{i,j}^{k}$ (9)
Next, we outline the learning and testing phases of CGCNs (See Appendix D for more details).
@@
4:As usual in kernel methods, we need to compute $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$, where $\mathbf{z}\in\mathbb{R}^{n\times F}$ is some input graph signal.
5:Form the vector $\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$ whose elements are kernel products (See Appendix D).
6:Compute the $i^{\text{th}}$ row of $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})$ using $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})_{i}=(Q^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$, where $(Q^{k})^{\dagger}$ is the pseudo-inverse of $Q^{k}$ (See Appendix D for more details).
@@
9:Layer $\ell$ induces a linear mapping between graph signals $\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}}\rightarrow\mathbb{R}^{n\times F_{\ell}}$ s.t. the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is $[\mathcal{X}_{\ell}]_{if}=[\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})]_{if}=\sum_{k=0}^{K}\langle\hat{\mathbf{z}}_{i}^{k}(\mathcal{X}_{\ell-1}),\hat{\mathbf{a}}_{\ell k}^{f}\rangle$, where $\hat{\mathbf{a}}_{\ell k}^{f}$ are the trainable parameters (filters) of the CGCN layer from (9).
@@ 4.3 The Algorithm for Learning Multi-Layer CGCNs
Algorithm 2 summarizes the training process of a multi-layer CGCN, estimating the parameters of each layer from bottom to top (i.e., in a layer-wise fashion). As such, to learn the nonlinear filter $\tau_{\ell}^{f}$ at layer $\ell$, we should learn the $P$-dimensional vector $\hat{\mathbf{a}}_{\ell k}^{f}$. For this sake, we define the graph signal $Z_{j}^{k}\in\mathbb{R}^{n\times P}$ for any $k\in[K]\cup\{0\}$ and $j\in[m]$ whose $i^{\text{th}}$ row is $[Q^{k}]_{(i,j)}$. We can then apply the nuclear norm relaxation from Section 4.1. Solving the resulting optimization problem yields a matrix of filters $\hat{\mathcal{A}}_{\ell k}\in\mathbb{R}^{P\times F_{\ell}}$ for layer $\ell$ whose $f^{\text{th}}$ column is $\hat{\mathbf{a}}_{\ell k}^{f}$.
@@
$\hat{\mathcal{A}}_{\ell}\in\operatorname*{arg\,min}_{\|\hat{\mathcal{A}}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}}\tilde{J}(\hat{\mathcal{A}}_{\ell}),\quad\text{ where }\tilde{J}(\hat{\mathcal{A}}_{\ell}):=\frac{1}{m}\sum_{j=1}^{m}J(\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1}^{(j)}),y^{(j)}).$ (10)
@@
where $\nabla_{\hat{\mathcal{A}}_{\ell}}\tilde{J}$ is the gradient of $\tilde{J}$ from (10) and $\Pi_{\mathcal{B}_{\ell}}$ denotes the Euclidean projection onto the nuclear norm ball $\{\hat{\mathcal{A}}_{\ell}:\|\hat{\mathcal{A}}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}\}$. This projection can be done by first computing the singular value decomposition of $\hat{\mathcal{A}}$ and then projecting the vector of singular values onto the $\ell_{1}$-ball via an efficient algorithm (see, e.g., [13, 12, 50]). The problem in (10) can also be solved using a projected version of any other optimizer (e.g., Adam [23]).
@@ 4.4 Scalability, Time Complexity and Challenges of Learning CGCNs
Algorithm 2’s runtime highly depends on how the factorization matrix $Q_{\ell}^{k}$ is computed and its width $P$ (line 3). Naively picking $P=nm$ requires solving the exact kernelized problem, which can be costly. To gain scalability, we compute a Nyström approximation [48] in $\mathcal{O}(P^{2}nm)$ time by randomly sampling training examples, computing their kernel matrix and representing each training example by its similarity to the sampled examples and kernel matrix. This avoids computing or storing the full kernel matrix, thus requiring only $\mathcal{O}(Pmn)$ space. While this approach is efficient [48, 9], it relies on the sampled subset adequately representing the entire data; poorly represented examples may degrade approximation quality. See Appendix E for more details and alternative methods.

### v2

@@ 4 CONVEXIFICATION OF GCNS
As $\mathcal{F}_{\text{gcn}}$ in (3) is a non-convex set and the activation function is nonlinear, the ERM problem on GCNs is non-convex, raising the challenges mentioned in Section 1. To overcome them, we propose a procedure for making both the ERM’s predictions domain and loss function convex. Following Zhang et al. (2017), we develop the class of Convexified GCNs (CGCNs). In Section 4.1, we first depict our method for linear activation functions. Given that the linear activation function would collapse all layers into one, this case is of no interest in practice. However, it supplies us with some insight into the general kernelization scheme of GCNs. In the nonlinear case (Section 4.2), we reduce the problem to the linear case by proposing a relaxation to a suitably chosen Reproducing Kernel Hilbert Space (RKHS), enabling us to obtain a convex formulation of the associated ERM problem. We then present an algorithm for learning CGCNs (Section 4.3), discuss its scalability and time complexity (Section 4.4), and establish its generalization guarantees (Section 4.5).
4.1 Linear Activation Functions
To provide the intuition for our method, we employ the nuclear norm in the simple case of the linear activation function $\sigma(\mathbf{w})=\mathbf{w}$, which provides us with a convex function class, unlike (3). In this case, the output of layer $\ell$ is simply $\mathcal{X}_{\ell}=\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}$ due to (1)-(2). Hence, $\mathcal{X}_{\ell}$ linearly depends on the filters $\{\mathcal{A}_{\ell k}\}_{k=0}^{K}$ of layer $\ell$, i.e., the convolution function $\Psi^{\mathcal{A}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}}\rightarrow\mathbb{R}^{n\times F_{\ell}}$ in (1) linearly depends on the parameters.
@@ 4.1 Linear Activation Functions
Nuclear Norm Relaxation. However, filters that satisfy the constraints (5a) and (5b) form a non-convex set. A standard convex relaxation of a rank constraint involves replacing it with a nuclear norm constraint. That is, controlling the sum of singular values instead of the number of non-zero ones, similarly to how the $\ell_{1}$-norm constraint can be viewed as a convex relaxation of the $\ell_{0}$-norm constraint.
Upper Bounding the Nuclear-Norm. In Appendix B, we prove that:
@@
4.2 Nonlinear Activation Functions
For certain nonlinear activation functions $\sigma$ and suitably chosen kernel function $\kappa$ such as the Gaussian RBF kernel, we next prove how the class $\mathcal{F}_{\text{gcn}}$ of GCNs in (3) can be relaxed to a Reproducing Kernel Hilbert Space (RKHS), reducing the problem to the linear activation case (Readers may refer to (Schölkopf et al., 2001) for a brief overview on RKHS). Beforehand, we rephrase the filter at layer $\ell$ from (1)-(2) as follows. Denoting the $f^{\text{th}}$ column of the filters $\mathcal{A}_{\ell k}$ as $\mathbf{a}^{f}_{\ell k}$, in Appendix D we prove that:
@@ 4.2 Nonlinear Activation Functions
A GCN induces, at each layer $\ell$, a nonlinear filter $\tau_{\ell}^{f}$ for the $f^{\text{th}}$ feature, which operates on the structured aggregation of neighborhood information and is given as (the activation $\sigma$ is applied entrywise):
@@
For any graph signal $\mathcal{S}^{k}\mathcal{X}$, let $\mathbf{z}_{i}^{k}(\mathcal{X}):=[\mathcal{S}^{k}\mathcal{X}]_{i}$ be the $i^{\text{th}}$ row of $\mathcal{S}^{k}\mathcal{X}$, which is the information aggregated at node $i$ from $k$-hops away. Defining the matrix $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k+1}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$, the output at layer $\ell$ is thus given by $\tau_{\ell}^{k}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$.
Kernelization of GCNs. In the sequel, we consider a training data $\mathcal{T}=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$, some layer of index $\ell$ and a positive semi-definite kernel function $\kappa_{\ell}:\mathbb{R}^{F_{\ell-1}}\times\mathbb{R}^{F_{\ell-1}}\rightarrow\mathbb{R}$ for layer $\ell$. Thus, for brevity, we omit indices involving $\ell$ and $\mathcal{X}_{\ell-1}^{(j)}$ when no ambiguity arise. Namely, we denote, e.g., $\mathcal{Z}_{i,j}:=\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}^{(j)})$ and $\mathbf{z}_{i,j}^{k}:=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})$.
First, we prove in Lemma 4.3 that the filter specified by (7) resides in the RKHS induced by the kernel (See Appendix E for a proof that adapts the one by Zhang et al. (2017) to GCNs). Zhang et al. (2017) prove that convolutional neural networks can be contained within the RKHS induced by properly chosen kernel and activation functions. For certain choices of kernels (e.g., the Gaussian RBF kernel) and a sufficiently smooth $\sigma$ (e.g., smoothed hinge loss), we can similarly prove that the filter $\tau_{\ell}^{f}$ in (7) is contained within the RKHS induced by the kernel function $\kappa_{\ell}$. See Appendices C.1-C.2 for possible choices of kernels and Appendix C.3 for valid activation functions, as well as an elaboration on sufficient smoothness in Remark C.2.
@@
Mercer’s theorem (Steinwart, 2008, Theorem 4.49) implies that the filter $\tau^{f}_{\ell}$ at layer $\ell$ for the $f^{\text{th}}$ feature can be reparametrized as a linear combination of kernel products, reducing the problem to the linear activations case. Formally:
@@
To avoid rederiving all results in the kernelized setting, we reduce the problem to the linear case in Section 4.1. Thus, consider the (symmetric) kernel matrix $\mathcal{K}^{k}_{\ell}\in\mathbb{R}^{nm\times nm}$, whose entry at row $(i,j)$ and column $(i^{\prime},j^{\prime})$ equals to $\kappa_{\ell}(\mathbf{z}_{i,j}^{k},\mathbf{z}_{i^{\prime},j^{\prime}}^{k})$. We factorize the kernel matrix as $\mathcal{K}^{k}_{\ell}=Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$ with $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ (as in (Zhang et al., 2017), we can use, e.g., Cholesky factorization (Dereniowski and Kubale, 2003) with $P=nm$).
@@
Thus, using Remark 4.5, the reparameterization in (8) of the filter at layer $\ell$ for the $f^{\text{th}}$ feature can be rephrased as:
$\tau^{f}_{\ell}(\mathcal{Z}_{i,j})$ $=\sum_{k=0}^{K}\langle\mathbf{q}_{i,j}^{k},\hat{\mathbf{a}}_{\ell k}^{f}\rangle;$ (9)
$\hat{\mathbf{a}}_{\ell k}^{f}:$ $=\sum_{(i^{\prime},j^{\prime},k)\in[n]\times[m]\times([K]\cup\{0\})}\alpha_{k,(i^{\prime},j^{\prime})}^{f}\mathbf{q}_{i^{\prime},j^{\prime}}^{k}$
Next, we outline the learning and testing phases of CGCNs (See Appendix E for more details).
@@
4:As usual in kernel methods, we need to compute $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$, where $\mathbf{z}\in\mathbb{R}^{F_{\ell-1}}$ is some input graph signal.
5:Form the vector $\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$ whose elements are kernel products (See Appendix E).
6:Compute the $i^{\text{th}}$ row of $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})$ using $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})_{i}=(Q^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$, where $(Q^{k})^{\dagger}$ is the pseudo-inverse of $Q^{k}$ (See Appendix E for more details).
@@
9:Layer $\ell$ induces a linear mapping between graph signals $\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}}\rightarrow\mathbb{R}^{n\times F_{\ell}}$ s.t. the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is $\mathcal{X}_{\ell}]_{if}=[\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})]_{if}=\sum_{k=0}^{K}\langle\hat{\mathbf{z}}_{i}^{k}(\mathcal{X}_{\ell-1}),\hat{\mathbf{a}}_{\ell k}^{f}\rangle$, where $\hat{\mathbf{a}}_{\ell k}^{f}$ are the trainable parameters (filters) of the CGCN layer from (9).
@@ 4.3 The Algorithm for Learning Multi-Layer CGCNs
Algorithm 2 summarizes the training process of a multi-layer CGCN, estimating the parameters of each layer from bottom to top (i.e., in a layer-wise fashion). As such, to learn the nonlinear filter $\tau_{\ell}^{f}$ at layer $\ell$, we should learn the $P$-dimensional vector $\hat{\mathbf{a}}_{\ell k}^{f}$. For this sake, we define the graph signal $Z_{j}^{k}\in\mathbb{R}^{n\times P}$ for any $k\in[K]\cup\{0\}$ and $j\in[m]$ whose $i^{\text{th}}$ row is $Q^{k}]_{(i,j)}$. We can then apply the nuclear norm relaxation from Section 4.1. Solving the resulting optimization problem yields a matrix of filters $\hat{\mathcal{A}}_{\ell k}\in\mathbb{R}^{P\times F_{\ell}}$ for layer $\ell$ whose $f^{\text{th}}$ column is $\hat{\mathbf{a}}_{\ell k}^{f}$.
@@
[…]
where $\tilde{J}(\hat{\mathcal{A}}_{\ell})=\frac{1}{m}\sum_{j=1}^{m}J(\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1}^{(j)}),y^{(j)}).$
@@
where $\nabla_{\hat{\mathcal{A}}_{\ell}}\tilde{J}$ is the gradient of $\tilde{J}$ from (10) and $\Pi_{\mathcal{B}_{\ell}}$ denotes the Euclidean projection onto the nuclear norm ball $\{\hat{\mathcal{A}}_{\ell}:\|\hat{\mathcal{A}}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}\}$. This projection can be done by first computing the singular value decomposition of $\hat{\mathcal{A}}$ and then projecting the vector of singular values onto the $\ell_{1}$-ball via an efficient algorithm (see, e.g., (Duchi et al., 2008, 2011; Xiao and Zhang, 2014)). The problem in (10) can also be solved using a projected version of any other optimizer (e.g., Adam (Kingma and Ba, 2015)).
@@ 4.4 Scalability, Time Complexity and Challenges of Learning CGCNs
Algorithm 2’s runtime highly depends on how the factorization matrix $Q_{\ell}^{k}$ is computed and its width $P$ (line 3). Naively picking $P=nm$ requires solving the exact kernelized problem, which can be costly. To gain scalability, we compute a Nyström approximation (Williams and Seeger, 2001) in $\mathcal{O}(P^{2}nm)$ time by randomly sampling training examples, computing their kernel matrix and representing each training example by its similarity to the sampled examples and kernel matrix. This avoids computing or storing the full kernel matrix, thus requiring only $\mathcal{O}(Pmn)$ space. While this approach is efficient (Williams and Seeger, 2001; Drineas and Mahoney, 2005), it relies on the sampled subset adequately representing the entire data; poorly represented examples may degrade approximation quality. See Appendix F for more details and alternative methods.